OpenStudy (anonymous):

A public library has fiction and nonfiction books in the ratio three colon two. If the library has three hundred sixty fiction books, how many nonfiction books does it have? seventy-two one hundred twenty one hundred eighty two hundred forty

OpenStudy (anonymous):

Just showed you this?

OpenStudy (anonymous):

fiction/ non fiction = 3/2 = 3x/2x 3x=360,x=120 non fiction = 2x=240

OpenStudy (anonymous):

:)